A proton is held at a fixed position. An electron is placed 0.95 m away and released. Find the electron's speed when it's 7.0 mm from the proton.
Answer #1

Conserving energies,

K.Ei + P.Ei = K.Ef + P.Ef ----------(1)

initial kinetic energy of electron K.Ei = 0

initial potential energy of electron   = P.E._i=\frac{1}{4\pi\epsilon_0}\frac{-(1.6\times10^{-19})^2}{0.95}

final kinetic energy of electron K.Ef = (1/2) mv2

final potential energy of electron   = P.E._f=\frac{1}{4\pi\epsilon_0}\frac{-(1.6\times10^{-19})^2}{7\times10^{-3}}

using equation (1) and mass of electron m = 9.1 x 10-31 kg

\frac{1}{4\pi\epsilon_0}\frac{-(1.6\times10^{-19})^2}{0.95}=\frac{1}{4\pi\epsilon_0}\frac{-(1.6\times10^{-19})^2}{7\times10^{-3}}+\frac{1}{2}mv^2

speed of electron v = 267.8 m/s